Consider the plane makes intercepts of 2 A , 3 A , 4 A on the coordinate axis of the crystal

with a: b: c = 4: 3 : 2. The crystal is orthorho1nbic crystal. Calculate the miller indices of

the given plane of the crystal.


1
Expert's answer
2021-10-11T16:45:21-0400

Intercept:

"\frac{2}{a}, \frac{3}{b}, \frac{4}{c} \\

= \frac{1}{2}, 1, 2"

Ratio of intercepts:

"\frac{2}{1}, 1, \frac{1}{2}"

Multiplying by 2 to convert into a lowest integer:

4, 2, 1

the miller indices of the given plane of the crystal:

(h,k,l) = (4, 2, 1)